# Quantum Neural Networks in Practice: A Comparative Study with Classical Models from Standard Data Sets to Industrial Images

### v1

@@ IV.4 Comparison of quantum model performances across data sets
With this explanation in mind, we may discuss Fig. 8 (a) and first observe that every model that performs well on any of the three data sets performs well on the second data set, which we attribute to the data set’s simplicity. The contrary, however, is not true. The models that perform well on the first and second data sets seem to perform rather mediocre on the third one and the best performing models of the third data set perform likewise mediocre on the first data set. We thus conclude that while our randomized generation of quantum neural networks seems to yield decent models that perform mostly good and at worst mediocre on all tested data sets, in order to have the best possible performance, customized models need to be crafted for each data set.
@@ V.2 Analysis of correlations between performance and hyperparameters
Figure 11: Correlation analysis for the quantum-classical hybrid convolutional neural networks (QCCNN) in terms of correlations between classification accuracies and various properties of the parametrized quantum circuit used for the quantum convolution. Each row corresponds to the results obtained for the data set indicated on the right. For simplicity and to have more points for each data set, we merge all QCCNNs and data set sizes used. Therefore, each panel contains results for models with all sizes of quantum-convolution filter, all entanglement structures of the circuits used for its quantum convolution, and number of layers in this quantum circuit, each trained with $N\in\{50,100,200,300,400,500\}$ data points. Visually, we still separate the latter two hyperparameters (entanglement structure and number of layers) by color. In panels (a) and (b), we analyze the correlation between classification accuracy and number of trainable parameters in the quantum circuit used in the quantum convolution. In panels (c) and (d), we analyze the correlation between classification accuracy and the number of layers in the quantum circuit. In panels (e) and (f), we analyze the correlation between classification accuracy and averaged gate entangling power $\bar{\mathcal{C}}$, cf. Eq. (18), of the models after training. Finally, panels (g) and (h) analyze the correlation between classification accuracy and the change of this average gate entanglement power before and after training, cf. Eq. (19). The QCCNNs without entangling gates are omitted in panels  (e)-(h). The linear lines in every panel — except panels (c) and (d) — show the general trend of all the data. In panels (c) and (d), the lines connect the average accuracy of each set of QCCNNs with a specific number of layers in its quantum circuit.
@@
Since the entangling gate structures were fixed and all entangling gates parametrized, the correlation between number of trainable parameters and number of entangling gates is very high. Therefore, we opt not to show this correlation plot. To analyze the usefulness of entanglement in QCCNN models, we calculate the averaged gate entangling power $\bar{\mathcal{C}}$, cf. Eq. (18), of the models after training, as done in Sec. IV.3. For this analysis, we exclude QCCNNs without entangling gates for obvious reasons. The relationship between this metric and classification accuracy is shown in Fig. 11 (e) and (f), for each data set. We do not find a correlation between the classification accuracy of a QCCNN and its averaged gate entangling power $\bar{\mathcal{C}}$. However, for the handwritten-digit data set, we observe that most models with larger $\bar{\mathcal{C}}$ reach a classification accuracy of $\sim 100\%$ on the validation data set. Conversely, if a model performs bad, it is most likely one with lower $\bar{\mathcal{C}}$. Nevertheless, the amount of data is again too limited to make a statistically significant statement. We still note that QCCNNs with more entangling gates show a lower $\bar{\mathcal{C}}$, which we further analyze in Fig. 12. In fact, these two quantities seem to be inversely proportional. Fitting the expression
@@
with $N_{2\mathrm{q}}$ the number of entangling gates in the quantum circuit used for the quantum convolution, to all data points, we get the curve depicted in Fig. 12, with parameters $a=1.47\pm 0.10$ and $b=0.13\pm 0.01$. We may explain this inverse relation as a consequence of the total gate entangling power being distributed among all entangling gates for each model. While a deeper analysis is needed, this suggests that, for problems which require a certain amount of entanglement generated by a quantum model, models with a small amount of entangling gates may be as performant as models with many such gates, while their training is easier to be done fully through classical simulation, since they can have much less parameters and, thus, require less circuit simulations to calculate gradients.
@@ VI Summary and conclusions
In this study, we use various data sets of increasing complexity — from simple, artificial data points up to complex, industry-level image data — to benchmark and compare the performance of classical and quantum machine learning methods in classifying these data sets. To do this, we consider various versions of each data set, in which we vary their sizes and consider either the full data or dimensionality-reduced versions of them. The latter is a prerequisite to make them accessible for computationally-intensive quantum machine learning models. We employ both convolutional autoencoders and principal component analyses as reduction methods for additional data variation. Moreover, we employ two different overall methodologies for classification. The first is the usage of randomized neural networks applied to the dimensionality-reduced data sets, while the second is the usage of convolutional neural networks applied to the original, non-reduced data sets. Fundamentally, both methodologies are comparisons between fully-classical and quantum-classical hybrid machine learning pipelines. While the first methodology applies either a fully-classical or fully-quantum neural network classifier to the dimensionality-reduced data, the inevitable dimensionality reduction as part of the total pipeline is always classical, which makes the methodology either fully-classical or quantum-classical hybrid. The same applies for the second methodology, where fully-quantum convolutional neural networks [31] would also, unfortunately, require too many resources. Thus, we resort to either fully-classical (CNNs) or quantum-classical hybrid convolutional neural networks (QCCNNs) [21].
Judging a model’s performance by the classification accuracy on the validation data set, we observe that the classical and quantum methods perform mostly similarly on almost all data set types and sizes. Our results from two simple data sets, one being the linear split of a $d$-dimensional hypercube and the other being 00’s and $1$’s from MNIST’s handwritten images [20], are in good agreement with previous studies [16], confirming that MNIST is trivial for both classical and quantum models while, for the linearly split hypercube, despite its simplicity, it is already surprisingly challenging to reach classification accuracies beyond $90\%$. However, these two simple data sets are merely used to benchmark our methodologies and validate its application for the third and most complex image data set from industry, which we focus on within our study. For this data set, we observe similar performance from classical and quantum models. Interestingly, we also find that the quantum and hybrid models, in comparison to fully-classical models, show a lower variance in their achieved classification accuracies with respect to initial training parameters, which may be regarded as advantageous for the former models when good initial training parameters are not known in advance. However, when analyzing the learned functions of classical and quantum models, we observe a tendency of the former to learn simple functions, while the latter have a tendency towards more complicated ones. While we do not analyze this in greater detail, it might hint at the absence of an inductive bias in the quantum models used [49].
Upon comparing our randomized quantum neural networks with some literature architectures, we observe that our best randomized quantum models can easily compete with the top performing literature model [39] among the ones we tried. Moreover, since we have benchmark results available for different data set types, we additionally use them to benchmark the performance of the best random quantum neural networks from one data set type on the other data set types. We observe that a top-performing model on any individual data set likely performs better than the average random model on the other data sets. However, designing top-performing models for each data set remains a data-set dependent task. When analyzing which hyperparameters make a quantum model perform well, we only find that having more trainable parameters available, in the case of quantum neural networks, seems to always help. Neither the type or structure of entangling gates, the feature encoding function, the measurement operator or the number of layers in the quantum neural networks or QCCNNs seem to have a statistically significant impact on the performance. Nevertheless, there are some interesting observations regarding entanglement. Due to our usage of trainable entangling gates, an analysis of whether gate entanglement power, based on gate concurrence, is turned up or down during a model’s training yields the following. For the random quantum neural networks, we observe no correlation between a model’s classification accuracy and the total averaged gate entanglement power or its change during training. In contrast, for the QCCNNs, while we do also not observe a correlation between classification accuracies and total averaged gate entanglement power, we in fact observe a positive change in averaged gate entanglement power due to training in virtually all scenarios. Independently, we also observe that the more entangling gates are available, the lower the per-gate entanglement power becomes. Hence, while there seems to be no trivial rule describing how gate entanglement power affects a model’s performance, its impact prompts further research. However, as an interesting observation, we find that most top-performing randomized quantum neural networks feature entangling gates, many of them even with an all-to-all structure. In contrast, for the quantum-classical hybrid neural networks, we observe that those model without entangling gates perform best in many of the scenarios.
@@
To conclude, we did compare the performance of fully-classical and quantum-classical hybrid machine learning models with the intent of judging their usefulness and applicability for an industrial data set. While we do not find that either of them outperforms the other in terms of achievable classification accuracy, there are certain subtleties that need to be taken into account. On the one hand, we needed to design the fully-classical model such as to allow a fair comparison in the first place. In detail, this implies, for instance, restricting the number of trainable parameters in the classical neural networks to be comparable to those of the random quantum neural networks. Hence, we did slightly hold back their full potential although a similar argument applies in favor of quantum models as well, which we kept classically simulable by restricting both qubit and gate count. On the other hand, the question of if one model outperforms another also strongly depends on the performance metric. Here, we exclusively use the classification accuracy on the validation data set, albeit other metrics such as convergence speed or training time could in principle be employed. However, for the latter, it would be difficult to make a tangible statement without running all quantum circuits on actual quantum hardware — instead of simulating them — and factor in data loading and circuit execution times as well as their frequent repetition until enough statistics would be gathered. Hence, until sufficiently potent quantum hardware is available, it is difficult to judge the practical potential of quantum machine learning methods, since, up to then, either a classical preprocessing of data or a restriction of quantum resources is required — both heavily impacting the conclusions one may draw from any results.
@@ Acknowledgements.
This work was supported by the German Federal Ministry for Education and Research (Bundesministerium für Bildung und Forschung) within the project ”AutoQML“ and project numbers 01MQ22002A and 01MQ22002B.
@@
D.B.: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing, Visualization. J.B.: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing, Visualization. C.T.: Funding Acquisition. F.S.: Conceptualization, Supervision, Funding Acquisition.

### v2

@@ IV.4 Comparison of quantum model performances across data sets
With this explanation in mind, we may discuss Fig. 9 (a) and first observe that every model which performs well on any of the three data sets also performs well on the second data set, which we attribute to the latter data set’s simplicity. The contrary, however, is not true. The models which perform well on the first and second data sets seem to perform rather mediocrely on the third one, and the best performing models on the third data set perform likewise mediocrely on the first data set. This separation seems to be correlated with the data’s complexity, cf. Fig. 1 (a)-(e), specifically with how separable the two classes are. While the first and second data sets are rather simple, in the sense that they are almost always linearly separable, the third data set is more complex and the classes are never linearly separable. We thus conclude that, while our randomized generation of quantum neural networks seems to yield decent models which perform mostly well, and at worst mediocrely on all tested data sets, in order to have the best possible performance, customized models need to be crafted in a data-set-dependent way.
@@ V.2 Analysis of correlations between performance and hyperparameters
Figure 12: Correlation analysis for the quantum-classical hybrid convolutional neural networks (QCCNN) in terms of correlations between classification accuracies and various properties of the parametrized quantum circuit used for the quantum convolution. Each row corresponds to the results obtained for the data set indicated on the right. For simplicity and to have more points for each data set, we merge all QCCNNs and data set sizes used. Therefore, each panel contains results for models with all sizes of quantum-convolution filter, all entanglement structures of the circuits used for its quantum convolution, and number of layers in this quantum circuit, each trained with $N\in\{50,100,200,300,400,500\}$ data points. Visually, we still separate the latter two hyperparameters (entanglement structure and number of layers) by color. In panels (a) and (b), we analyze the correlation between classification accuracy and number of trainable parameters in the quantum circuit used in the quantum convolution. In panels (c) and (d), we analyze the correlation between classification accuracy and the number of layers in the quantum circuit. In panels (e) and (f), we analyze the correlation between classification accuracy and averaged gate entangling power $\bar{\mathcal{C}}$, cf. Eq. (18), of the models after training. Finally, panels (g) and (h) analyze the correlation between classification accuracy and the change of this average gate entanglement power before and after training, cf. Eq. (19). The QCCNNs without entangling gates are omitted in panels  (e)-(h). The linear lines in every panel — except panels (c) and (d) — solely show the general trend of all the data and should not be taken as a fit to the data. In panels (c) and (d), the lines connect the average accuracy of each set of QCCNNs with a specific number of layers in its quantum circuit.
@@
Since the entangling gate structures were fixed and all entangling gates parametrized, the correlation between number of trainable parameters and number of entangling gates is very high, by definition. Therefore, we opt not to show this correlation plot. To analyze the usefulness of entanglement in QCCNN models, we calculate the averaged gate entangling power $\bar{\mathcal{C}}$, cf. Eq. (18), of the models after training, as done in Sec. IV.3. For this analysis, we exclude QCCNNs without entangling gates for obvious reasons. The relationship between this metric and classification accuracy is shown in Fig. 12 (e) and (f), for each data set. We do not find a correlation between the classification accuracy of a QCCNN and its averaged gate entangling power $\bar{\mathcal{C}}$. However, for the handwritten-digit data set, we observe that most models with larger $\bar{\mathcal{C}}$ reach a classification accuracy of $\sim 100\%$ on the validation data set. Conversely, if a model performs bad, it is most likely one with lower $\bar{\mathcal{C}}$. Nevertheless, the amount of data is again too limited to make a statistically significant statement. We still note that QCCNNs with more entangling gates show a lower $\bar{\mathcal{C}}$, which we further analyze in Fig. 13. In fact, these two quantities seem to be inversely proportional. Fitting the expression
@@
with $N_{2\mathrm{q}}$ the number of entangling gates in the quantum circuit used for the quantum convolution, to all data points, we get the curve depicted in Fig. 13, with parameters $a=1.47\pm 0.10$ and $b=0.13\pm 0.01$. We may explain this inverse relation as a consequence of the total gate entangling power being distributed among all entangling gates for each model. While a deeper analysis is needed, this suggests that, for problems which require a certain amount of entanglement generated by a quantum model, models with a small amount of entangling gates may be as performant as models with many such gates. Moreover, training the former should be easier through classical simulation, since they can have much less parameters and, thus, require less circuit simulations to calculate gradients.
@@ VI Summary and conclusions
In this study, we conduct a systematic investigation of classical and quantum machine learning approaches for binary image classification across data sets of increasing complexity. Our work provides a uniquely comprehensive hyperparameter optimization analysis — rarely conducted in quantum machine learning literature — offering insights into which quantum circuit properties actually contribute to model performance. We evaluate both classical and quantum-classical hybrid approaches using three progressively complex data sets: (i) artificial hypercube data, (ii) MNIST handwritten digits (0’s and 1’s) [23], and (iii) real-world industrial images from laser cutting machines.
Our methodological framework employs two complementary approaches: (1) randomized (quantum) neural networks applied to dimensionality-reduced data, and (2) (hybrid) convolutional neural networks applied to full image data. Both methodologies effectively compare fully-classical versus quantum-classical hybrid pipelines. Dimensionality reduction through convolutional autoencoders (CAE) and principal component analysis (PCA) serves as a necessary preprocessing step for fully-quantum models, as the high dimensionality of industrial data sets would otherwise make them computationally intractable for quantum processing. The first methodology employs either classical or quantum neural network classifiers on dimensionality-reduced data, while the second uses either classical CNNs or quantum-classical hybrid convolutional neural networks (QCCNNs) [24].
Quantitatively, using classification accuracy on validation data as our performance metric, we observe that classical and quantum methods achieve statistically equivalent classification accuracies across most tested data sets. For the industrial image data, maximum averaged accuracies of quantum models reached approximately $80\%$ for dimensionality-reduced data with $d=8$ features and between $95\%$ and $100\%$ for full-image classification with convolutional approaches, except for the smallest data sets containing 50 and 100 total images. Our results from the two simple data sets, the linear split of a $d$-dimensional hypercube and the 00’s and $1$’s from MNIST’s handwritten images, are in good agreement with previous studies [19], confirming that MNIST is trivial for both classical and quantum models while, for the linearly split hypercube, despite its simplicity, it is already surprisingly challenging to reach classification accuracies beyond $90\%$. However, these two simple data sets are merely used to benchmark our methodologies and validate its application for the industry data set. This empirical equivalence in performance highlights a significant finding: despite substantial theoretical advancement in quantum machine learning, practical implementation on realistic industrial data sets shows no clear quantum advantage in classification accuracy.
An intriguing finding is that quantum and hybrid models demonstrate significantly lower variance in classification accuracies with respect to initial training parameters compared to classical approaches. This was consistently observed across all data sets and methodologies, suggesting improved stability in quantum models when optimal initialization parameters are unknown — a potentially valuable property for practical applications. Conversely, when analyzing learned decision boundaries, classical models consistently produced simpler functions while quantum models generated notably more complex classification boundaries, even for data sets with easily-separable classes. This tendency toward complexity despite no performance advantage might indicate a lack of appropriate inductive bias in the quantum models used [52].
Our comprehensive hyperparameter analysis revealed surprisingly few consistent correlations between model architecture choices and performance. Among all parameters analyzed (feature encoding methods, number of circuit layers, entangling gate type and structure, gate entangling power, measurement operators), only the number of trainable parameters showed consistent positive correlation with model performance. However, this is not a general statement as, at some point, we expect to run into the overfitting regime. The inconsistent importance of hyperparameters across our experiments — particularly regarding entanglement structure and gate entangling power — illuminates the current limited theoretical understanding of how quantum models actually function and what determines their performance.
Upon comparing our randomized quantum neural networks with some literature architectures, we observe that our best randomized quantum models can easily compete with the top performing literature model [2] among the ones we tried, including Chebyshev [1], Hubregtsen [2], and HighDim [3]. For the randomized quantum neural networks, we found no correlation between model performance and total averaged gate entanglement power or its change during training, with gate entangling power quantified using gate concurrence [54]. However, $\sim 94\%$ of the best performing models did feature entangling gates, often with all-to-all connectivity. However, we note that this entanglement structure can always be mapped to a linear one with a polynomial gate overhead [55]. Thus, while the type of entangling structure may not be a crucial factor, the non-negligible entangling power of the top-performing models is a hopeful observation. Further testing the usefulness of entanglement in quantum circuits with a fixed amount of trainable parameters is a goal for future work.
For QCCNNs, models without entangling gates performed equally well as those with entanglement in most scenarios, though they required more iterations to converge in both classification accuracy and cross-entropy loss, especially with limited training data. This suggests that, while entanglement may not be a crucial factor for the performance of quantum-classical hybrid models, it may still be beneficial for their training. Interestingly, for the QCCNNs, we observe a positive change in averaged gate entanglement power due to training in virtually all scenarios. We also observed an inverse relationship between the number of entangling gates and per-gate entanglement power, suggesting that the total entanglement capability is distributed among available gates — a finding with implications for efficient quantum circuit design. However, a deeper statistical analysis would be needed to confirm the generality of these properties.
Our cross-dataset analysis reveals that quantum-model transferability is limited, with performance possibly dependent on data separability characteristics. Models performing well on linearly-separable data showed mediocre performance on the more complex, non-linearly separable industrial images. While there seems to be a correlation between model architecture and class separability, further research is needed to confirm this hypothesis. This suggests that despite the universal approximation capabilities theoretically attributed to quantum models, custom architecture design remains necessary for optimal performance on specific data sets.
We approached this study with fairness considerations in both directions: restricting the parameter count of classical models to match quantum models, while limiting quantum model size to ensure classical simulability. Despite these constraints, our findings provide valuable insights into the current state of quantum machine learning for industrial image classification. The absence of performance advantage, combined with our systematic hyperparameter optimization analysis, illustrates the significant gap between theoretical quantum advantage propositions and practical implementations.
@@
The question of whether one model outperforms another also strongly depends on the performance metric. Here, we exclusively use the classification accuracy on the validation data set, although other metrics such as convergence speed or training time could, in principle, be employed. However, for the latter, it would be difficult to make a tangible statement without running all quantum circuits on actual quantum hardware — instead of simulating them — and factor in data loading and circuit execution times as well as their frequent repetition until enough statistics would be gathered. Hence, until sufficiently potent quantum hardware is available, it is difficult to judge the practical potential of quantum machine learning methods, since, up to then, either a classical preprocessing of data or a restriction of quantum resources is required — both heavily impacting the conclusions one may draw from any results. Nevertheless, our comprehensive hyperparameter study provides a methodological framework and empirical baseline for future quantum machine learning research focused on realistic industrial applications, offering crucial insights into the current limitations and potential directions for advancement in this rapidly evolving field.
@@ Acknowledgements.
This work was supported by the Federal Ministry for Economic Affairs and Climate Action (Bundesministerium für Wirtschaft und Klimaschutz) within the project ”AutoQML“ and project numbers 01MQ22002A and 01MQ22002B.
@@
D.B.: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing, Visualization. J.B.: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing, Visualization. C.T.: Funding Acquisition, Supervision. F.S.: Conceptualization, Supervision, Funding Acquisition.